1. 2.50 L of 0.450 mol/L solution of silver nitrate is mixed with a solution prepared by dissolving 86.2 g sodium phosphate.
A) Which reactant is the limiting Reagent? Show work
B) what chemical amount (moles) of the excess will remain?
C) what mass of precipitate will be produced?

Unit - Stoichiometry
1
Expert's answer
2020-12-17T07:40:49-0500

3AgNO3 + Na3PO4 → Ag3PO4 + 3NaNO3

A) n(AgNO3) "= 2.5 \times 0.45 = 1.125 \;mol"

M(Na3PO4) = 163.94 g/mol

n(Na3PO4) "= \frac{86.2}{163.94} = 0.525 \;mol"

According to the reaction:

n(AgNO3) : n(Na3PO4) = 3 : 1

Real quantities:

n(AgNO3) : n(Na3PO4) = 1.125 : 0.525 = 2.14 : 1

So, silver nitrate is a limiting reagent.

B) n_1(Na3PO4) "= \frac{1}{3}n(AgNO_3) = \frac{1}{3} \times 1.125 = 0.375 \;mol" (used in reaction)

Δn(Na3PO4) = 0.525 – 0.375 = 0.15 mol (the excess will remain)

C) n(Ag3PO4) "= \frac{1}{3}n(AgNO3) = \frac{1}{3} \times 1.125 = 0.375 \;mol"

M(Ag3PO4 ) = 418.58 g/mol

m(Ag3PO4) "= 0.375 \times 418.58 = 156.96 \;g"